# Thermodynamic analysis and shadow bound of black holes surrounded by a dark matter halo

### v1

@@ 4 Shadow bounds
$r_{\rm RL}(M_{b},M,a_{0})$ (21)
whose expression is too complicated to write down here. For $M_{b}\ll M\ll a_{0}$, however, one obtains an approximate light ring
$r_{\rm{RL},a}=3M_{\rm bh}\Big{(}1+\frac{M_{b}M}{a_{0}^{2}}\Big{)}$ (22)
@@
Figure 5: Plot of the exact critical impact parameter $b_{c}(M_{b}=1,M,a_{0})$ as functions of $M\in[1,1000]$ and $a_{0}\in[0,5000]$ with $M_{b}=1$. One has the lower limit of $b_{c}(1,0.001,5000)=b_{c}^{S}=5.196$. Different black lines denote the lines for the same critical impact parameters. Based on the EHT results, the favored region ($5.196\leq b_{c}\leq 5.716$) is separated from the disfavored region ($b_{c}>5.716$) by the cyan line ($b_{c}^{B}=5.716$). This line could be represented precisely by $a_{0}(M)=10.8M$. For the SgrA, the magenta line ($b_{c}^{K}=5.5$) is given by $a_{0}(M)=18M$, while the blue line ($b_{c}^{V}=5.3$) is given by $a_{0}(M)=50M$.
@@
Three types of critical impact parameters are given by
@@
where the last is obtained by making large $a_{0}$-approximation on $b_{c}$ [25]. The middle corresponds to the Newtonian approximation which is consistent with thermodynamic analysis. These approximate parameters are similar to those of heat capacity except replacing $r=r_{\rm RL}$ by $r=r_{+}$. We check that $\lim_{M\to 0}b_{c}=3\sqrt{3}M_{b}(=b_{c}^{S}$, Schwarzschild case). In Fig. 5, we show the favored (disfavored) region by making use of the results of the EHT collaboration for M87 [1] and for SgrA [6]. Here, we have to use the exact critical impact parameter $b_{c}$ only. The lower limit is found to be $b_{c}^{L}(1,0.001,5000)=b_{c}^{S}=5.196$. If a relative deviation from the Schwarzschild result ($b_{c}^{S}$) is less (greater) than 10%, the solution is in the favored (disfavored) region [23]. The line of $b_{c}^{E}=5.716$ represents the upper limit for the favored region and it is also given by $a_{0}(M)=10.8M~{}({\cal C}=1/10.8)$. Therefore, the favored region for shadow bound is represented by $a_{0}\geq 10.8M$. For the SgrA* [6], the favored region is represented by $a_{0}\geq 18M(b_{c}^{K}=5.5,~{}{\cal C}=1/18)$ for Keck Observatory and it is small as $a_{0}\geq 50M(b_{c}^{V}=5.3,~{}{\cal C}=1/50)$ for the Very Large Telescope Interferometer. We note that the observation of galaxies corresponds to the regime of $a_{0}\geq 10^{4}M$ [27].
@@
Furthermore, we need to introduce the favored region when adopting two approximated critical impact parameters $b_{c}^{N}$ and $b_{c}^{a_{0}^{-2}}$. If the relative deviation from the Schwarzschild result is less than 10%, the solution is in the favorable region with [5.196(E--),5.716(E+)] for M87. Also, we introduce the favored region [4.5(K--),5.5(K+)] for Keck Observatory of SgrA. As is shown in (Left) Fig. 6, three impact parameters $b_{c}^{N}(1,M,5000),$ $b_{c}^{a_{0}^{-2}}(1,M,5000),b_{c}(1,M,5000)$ belong to the favored region starting from $M=0$ to $M=500,465,462$ for M87 and $292,280,279$ for SgrA, respectively. Here $b_{c}^{N}$ selects the largest $M$. From (Right) Fig. 5, one finds that all of $b_{c}^{N}(M_{b},10,5000),b_{c}^{a_{0}^{-2}},b_{c}$ are in the favored region from $M_{b}=0.999$ to 1.098 for M87 and from $M_{b}=0.864$ to 1.056 for SgrA.
@@ 5 Discussions
We have investigated thermodynamics and shadow bound of a BH immersed in a DMH (combined geometry). The redshift factor $e^{\Upsilon(r)}$, representing a key feature of the BH surrounded by DMH, has arisen from the mass function of DMH distributed over outside the event horizon. First of all, analyzing the heat capacity of the BH surrounded by DMH has indicated that the BH could not be in thermal equilibrium with the DMH in any regions, contradicting to the Schwarzschild-AdS BH [28, 29]. This means that the thermodynamic influence of the environment (DMH) is relatively small on the BH and it does not alter the nature of the BH with negative heat capacity, being consistent with its influence of the DMH on classical and quantum radiation around the BH [12]. One may need a cavity enclosing the combined geometry to have a positive heat capacity.
The Newtonian ($1/a_{0}$) approximation gave use a correct thermodynamic description for the BH surrounded by DMH because the first law of thermodynamics and Smarr formula are satisfied. On the other hand, the $1/a^{2}_{0}$ approximation is not suitable for describing thermodynamics of the BH surrounded by DMH because its Smarr formula is violated. We prefer the Newtonian approximation than the $1/a^{2}_{0}$ approximation. We have used the Newtonian Helmholtz free energy to show that there is no phase transition to other BH with positive heat capacity surrounded by a DMH.
We have used the M87* shadow data obtained by the EHT collaboration and SgrA data from Keck observatory and Very Large Telescope Interferometer to constrain two parameters ($M,a_{0}$) of DMH surrounding a BH with mass $M_{b}=1$. The shadow cast is usually described by the exact critical impact parameter $b_{c}$. We accept that if the relative deviation from the Schwarzschild result ($b_{c}^{S}=3\sqrt{3}M_{b}=5.196$) is less (greater) than 10%, the solution is in the favorable (disfavored) region [23]. The line of $b_{c}^{E}=5.716$ represents the upper limit and it is also given by $a_{0}(M)=10.8M$. Therefore, the shadow bound for favored region is represented by $a_{0}\geq 10.8M$. For the SgrA* [6], the shadow bound is represented by $a_{0}\geq 18M(b_{c}^{K}=5.5)$ for Keck Observatory and $a_{0}\geq 50M(b_{c}^{V}=5.3)$ for the Very Large Telescope Interferometer. In addition, the favored regions were set when including two approximated critical impact parameters $b_{c}^{N}(1,M,5000)$ and $b_{c}^{a_{0}^{-2}}(1,M,5000)$.

### v2

@@ 4 Shadow bounds
$r_{\rm RL}(M_{b},M,a_{0})$ $=$ $M+M_{b}-\frac{2a_{0}}{3}$ (21)
- $\frac{2^{1/3}\zeta}{3(\eta+\sqrt{4(\zeta^{3}+\eta^{2})})^{1/3}+\frac{1}{3\cdot 2^{1/3}}(\eta+\sqrt{4(\zeta^{3}+\eta^{2})})^{1/3}}$
with
$\zeta$ $=$ $-a_{0}^{2}+12a_{0}M-9(M^{2}+M_{b}^{2})-6a_{0}M_{b}+18MM_{b},$ (22)
$\eta$ $=$ $2a_{0}^{3}+45a_{0}^{2}M-108a_{0}M^{2}+54(M^{3}+a_{0}M_{b}^{2}+M_{b}^{3})$ (23)
$+$ $18a_{0}^{2}M_{b}+162(a_{0}MM_{b}-M^{2}M_{b}).$
For $M_{b}\ll M\ll a_{0}$, however, one obtains an approximate light ring
$r_{\rm{RL},a}=3M_{b}\Big(1+\frac{M_{b}M}{a_{0}^{2}}\Big)$ (24)
@@
Figure 5: Plot of the exact critical impact parameter $b_{c}(M_{b}=1,M,a_{0})$ as functions of $M\in[1,1000]$ and $a_{0}\in[0,5000]$ with $M_{b}=1$. One has the lower limit of $b_{c}(1,0.001,5000)=b_{c}^{S}=5.196$. Different black lines denote the lines for the same critical impact parameters. Based on the EHT results, the favored (consistent with observations) region ($5.196\leq b_{c}\leq 5.716$) is separated from the disfavored (inconsistent with observations) region ($b_{c}>5.716$) by the cyan line ($b_{c}^{E}=5.716$). This line could be represented precisely by $a_{0}(M)=10.8M$. For the SgrA, the magenta line ($b_{c}^{K}=5.5$) is given by $a_{0}(M)=18M$, while the blue line ($b_{c}^{V}=5.3$) is given by $a_{0}(M)=50M$.
@@
Three types of critical impact parameters are available as
@@
where the last is obtained by making large $a_{0}$-approximation on $b_{c}$ [35]. The middle corresponds to the Newtonian approximation which is consistent with thermodynamic analysis. This approximation came from $e^{\Upsilon(r)}=1-2M/a_{0}$ in the redshift factor when closing to the central BH. In this approximation, there is no correction to the light ring as is shown in Eq.(24). Here, the DMH distribution ($M,a_{0}$) outside the BH determines a modification of the redshift factor and critical impact parameter solely, as was found in Newtonian boson star profiles [46, 47]. We find from Eq.(28) that the leading order correction is due to gravitational redshift, but there are subdominant contributions whose detection is challenging if $M/a_{0}\leq 10^{-4}$. It shows that the galactic content can affect the shadow of BH by terms of order $M^{2}/a_{0}^{2}\leq 10^{-8}$.
We check that $\lim_{M\to 0}b_{c}=3\sqrt{3}M_{b}(=b_{c}^{S}$, Schwarzschild case). In Fig. 5, we show the favored (disfavored) region by making use of the results of the EHT collaboration for M87 [11] and for SgrA [16]. Here, we have to use the exact critical impact parameter $b_{c}$ only. The lower limit is found to be $b_{c}^{L}(1,0.001,5000)=b_{c}^{S}=5.196$. If a relative deviation from the Schwarzschild result ($b_{c}^{S}$) is less (greater) than 10%, the solution is in the favored (disfavored) region [33]. The line of $b_{c}^{E}=5.716$ represents the upper limit for the favored region and it is also given by $a_{0}(M)=10.8M~({\cal C}=1/10.8)$. Therefore, the favored region for shadow bound is represented by $a_{0}\geq 10.8M$. For the SgrA* [16], the favored region is represented by $a_{0}\geq 18M(b_{c}^{K}=5.5,~{\cal C}=1/18)$ for Keck Observatory and it is small as $a_{0}\geq 50M(b_{c}^{V}=5.3,~{\cal C}=1/50)$ for the Very Large Telescope Interferometer. However, we remind the reader that the observation of galaxies corresponds to the regime of $a_{0}\geq 10^{4}M$ [37].
@@
Furthermore, we need to introduce the favored region when adopting two approximated critical impact parameters $b_{c}^{N}$ and $b_{c}^{a_{0}^{-2}}$. If the relative deviation from the Schwarzschild result is less than 10%, the solution is in the favorable region with [5.196(E-),5.716(E+)] for the EHT observations (M87). Also, we introduce the favored region [4.5(K-),5.5(K+)] for Keck Observatory of SgrA. As is shown in (Left) Fig. 6, three impact parameters $b_{c}^{N}(1,M,5000),$ $b_{c}^{a_{0}^{-2}}(1,M,5000),b_{c}(1,M,5000)$ belong to the favored region starting from $M=0$ to $M=500,465,462$ for M87 and $292,280,279$ for SgrA, respectively. Here $b_{c}^{N}$ selects the largest $M$. From (Right) Fig. 5, one finds that all of $b_{c}^{N}(M_{b},10,5000),b_{c}^{a_{0}^{-2}},b_{c}$ are in the favored region from $M_{b}=0.999$ to 1.098 for M87 and from $M_{b}=0.864$ to 1.056 for SgrA.
@@ 5 Discussions
We have investigated thermodynamics and shadow bound of the BH immersed in the DMH. In this case, we wish to point out that the redshift factor $e^{\Upsilon(r)}$ has captured a key feature of the DMH surrounding the BH, while the BH is still described by Schwarzschild BH. This factor has arisen from the mass function of DMH which describes the mass distribution over outside the event horizon.
First of all, studying the heat capacity of the BH surrounded by DMH implied that the BH could not be in thermal equilibrium with the DMH in any regions. This contradicts to the Schwarzschild-AdS BH [38, 39]. This means that the thermodynamic influence of the DMH (environment) is limited to being small on the BH. Therefore, it does not alter the nature of the BH with negative heat capacity which implies that the BH is thermodynamically unstable. This phenomena is similar to the influence of the DMH on classical and quantum radiation around the BH [22]. One may introduce either cavity or AdS spacetime enclosing the combined geometry to obtain a thermodynamically stable BH.
In addition, it is desirable to note that the Newtonian ($1/a_{0}$) approximation provides a correct thermodynamic description for the BH surrounded by DMH because the first law of thermodynamics and Smarr formula are satisfied. On the other hand, the $1/a^{2}_{0}$ approximation is inappropriate for describing its thermodynamics because its Smarr formula is violated. In this respect, we have employed the Newtonian Helmholtz free energy to show that there is no phase transition to other BHs with positive heat capacity.
We have used the M87* and SgrA shadow data to constrain a relation between two parameters ($M,a_{0}$) of the DMH surrounding the central BH with mass $M_{b}=1$. The shadow radius was analyzed by the exact critical impact parameter $b_{c}$. If a relative deviation from the Schwarzschild result ($b_{c}^{S}=3\sqrt{3}M_{b}=5.196$) is less (greater) than 10%, the solution is in the favorable (disfavored) region [33]. Here, the line of $b_{c}^{E}=5.716$ represents the upper limit and it corresponds to $a_{0}(M)=10.8M(\mathcal{C}=1/10.8)$ [34]. Therefore, the shadow bound for favored (consistent) region could be represented by $a_{0}\geq 10.8M$.
It is desirable to note that the authors in [35] have predicted the compactness of DMH ($\mathcal{C}=1/12\sim 1/15$), while the authors [33] have proposed an interesting case of $\mathcal{C}=3/2$. On the other hand, the shadow bound is determined by $a_{0}\geq 18M(b_{c}^{K}=5.5,~\mathcal{C}=1/18)$ for Keck Observatory and $a_{0}\geq 50M(b_{c}^{V}=5.3,~\mathcal{C}=1/50)$ for the Very Large Telescope Interferometer [16]. However, these $\mathcal{C}$ all are greater than observations of galaxy $a_{0}\geq 10^{4}M(b_{c}^{V}=5.3)\to\mathcal{C}=10^{-4}]$. In the future, one has to explain this big difference if the present model is suitable for describing the BH surrounded by the DMH.